# The Bisognano-Wichmann property for non-unitary Wightman conformal field theories

### v1

@@ 2.2 Wightman CFTs and vertex algebras
Our main object of study is potentially non-unitary Möbius-covariant Wightman conformal field theories on the unit circle $S^{1}\subset{\mathbb{C}}$. In this context, the Möbius group $\operatorname{M\ddot{o}b}$ refers to $\operatorname{PSU}(1,1)$, the group of holomorphic automorphisms of the unit disk. Such theories were systematically studied in [CRTT24] (see also [RTT22]). In contrast to Section 2.1, we do not limit ourselves to unitary theories (which possess a compatible inner product). We will summarize here the basic notions, and the reader may consult the reference for further details.
@@
We use the Fréchet topology induced on $C^{\infty}(S^{1})$ by the $C^{N}$ norms [^4] In [RTT22, CRTT24] we used the Sobolev norms $\left\|f\right\|_{H^{N}}$, but the resulting topologies are equivalent due to embeddings $H^{N}(S^{1})\subset C^{N^{\prime}}(S^{1})$ and $C^{M}(S^{1})\subset H^{M^{\prime}}(S^{1})$.
$\left\|f\right\|_{C^{N}}=\sum_{j=1}^{N}\big{\|}f^{(j)}\big{\|}_{\infty}=\sum_{j=1}^{N}\big{\|}\tfrac{d^{j}}{d\theta^{j}}f\big{\|}_{\infty}.$
@@
The Wightman axioms have not been extensively studied in their own right within the context of axiomatic 2d chiral conformal field theory, but it turns out that they are exactly equivalent to Möbius vertex algebras, along with a chosen set of quasiprimary vectors which generate the vertex algebra. It has been long-known that this correspondence is “morally valid” (see e.g. [Kac98, §1.2]), certainly in the case of unitary theories, but it was made precise in [CRTT24]. In the case of vertex algebras, the Möbius covariance condition is with respect to the complexified Lie algebra $\operatorname{Lie}(\operatorname{M\ddot{o}b})_{\mathbb{C}}$, which is spanned by elements $\{L_{-1},L_{0},L_{1}\}$ satisfying the familiar relation $[L_{n},L_{m}]=(n-m)L_{n+m}$.
@@
    $\Omega$ is $\operatorname{Lie}(\operatorname{M\ddot{o}b})$-invariant, i.e. $L_{m}\Omega=0$ for $m=-1,0,1$.
@@
    $[L_{m},Y(v,z)]=\sum_{j=0}^{m+1}\binom{m+1}{j}z^{m+1-j}Y(L_{j-1}v,z)$ and $Y(L_{-1}v,z)=\frac{d}{dz}Y(v,z)$ for all $v\in{\mathcal{V}}$ and $m=-1,0,1$.
@@
In the other direction, given a Möbius vertex algebra ${\mathcal{V}}$ with a generating set of quasiprimary vectors, there is a canonical extension ${\mathcal{D}}\supset{\mathcal{V}}$ (which can be realized as a subspace of the algebraic completion of ${\mathcal{V}}$) and Möbius-covariant Wightman CFT on ${\mathcal{D}}$ corresponding to ${\mathcal{V}}$ in the sense described above. This correspondence can be made into an equivalence of categories [CRTT24, §3.3].
@@
In both settings, we have a notion of an invariant Hermitian form, which we call an involutive theory in general, and a unitary theory if the Hermitian form is an inner product. Some key examples of involutive but non-unitary theories include the non-unitary minimal models, such as the Yang-Lee model. These notions will be outlined in detail in Sections 4 and 6, respectively. Invariant forms for a Wightman theory exactly correspond to invariant forms for the corresponding vertex algebra [CRTT24, §4].
We close with a brief preview of the themes of the article. Following [SW64], given a Wightman CFT ${\mathcal{F}}$ acting on its domain ${\mathcal{D}}$, for every interval [^5] The term ‘interval’ in $S^{1}$ will be used exclusively for open, nonempty, proper, non-dense connected subsets of $S^{1}$. $I\subset S^{1}$ we form the polynomial algebra ${\mathcal{P}}(I)\subset\operatorname{End}({\mathcal{D}})$, which is to say the unital algebra generated by operators $\varphi(f)$ with $\operatorname{supp}(f)\subset I$. We obtain a net ${\mathcal{P}}$ of algebras which formally satisfies many of the properties of a Haag-Kastler net, but is in fact encoding the data of a Möbius vertex algebra. From [CRTT24, App. A], this net has the Reeh-Schlieder property, which is to say that ${\mathcal{P}}(I)\Omega$ is (${\mathcal{F}}$-strong) dense in ${\mathcal{D}}$ for every interval $I$, and if $x\in{\mathcal{P}}(I)$ and $x\Omega=0$, then $x=0$. Continuing along the same line, we have the following motivating questions:
@@ 3 Analytic continuation for non-unitary Wightman CFTs
For an interval $I\subset S^{1}$ we let ${\mathcal{P}}(I)\subset\operatorname{End}({\mathcal{D}})$ be the unital algebra generated by operators $\varphi(f)$ with $\varphi\in{\mathcal{F}}$ and $\operatorname{supp}f\subset I$. For $\tau\in{\mathbb{C}}$, let ${\mathbb{S}}_{\tau}$ be the closed strip bounded by the lines ${\mathbb{R}}$ and ${\mathbb{R}}+\tau$. Our first goal is to investigate the analytic continuation of $V_{t}$ to the closed strip
[…]
@@
$(f_{1},\ldots,f_{k})\mapsto\lambda\big{(}\varphi_{1}(f_{1})\cdots\varphi_{k}(f_{k})\Omega\big{)}$
are (jointly) continuous in the $f_{j}\in C^{\infty}(S^{1})$.
@@
By the Hahn-Banach theorem such a vector $\Psi$ is uniquely determined.
If $\tau=t\in{\mathbb{R}}$ then $D(\widetilde{V}_{t})={\mathcal{D}}$ and $\widetilde{V}_{t}=V_{t}$. The tilde in the notation $\widetilde{V}_{\tau}$ and $D(\widetilde{V}_{\tau})$ is present to distinguish from the domain $D(V_{\tau})$ considered in Section 5 in the context of CFTs with an invariant Hermitian form. If $\tau\in{\mathbb{C}}$ and $t\in{\mathbb{R}}$ we have $\widetilde{V}_{\tau+t}=\widetilde{V}_{\tau}V_{t}=V_{t}\widetilde{V}_{\tau}$ as partially defined operators [^7] This means that the operators in question have the same domain, where the domain of the composite of partially defined operators $X_{1}X_{2}$ consists of all vectors $\Phi$ such that $\Phi\in D(X_{2})$ and $X_{2}\Phi\in D(X_{1})$.. More generally for $\tau_{1},\tau_{2}\in{\mathbb{C}}$ the operator $V_{\tau_{1}+\tau_{2}}$ is an extension of $V_{\tau_{1}}V_{\tau_{2}}$. It is clear that $\Omega\in D(\widetilde{V}_{\tau})$, but it is not immediately obvious that $D(\widetilde{V}_{\tau})$ is any larger. Thus our first step is to identify a ‘large’ subspace of $D(\widetilde{V}_{\tau})$.
Let $C_{0}^{\infty}(I_{+})$ and $C_{0}^{\infty}(I_{-})$ be smooth functions on $I_{+}$ (resp. $I_{-}$) that vanish to all orders at the endpoints. We will sometimes embed these spaces into $C^{\infty}(S^{1})$ by extending with the value 00, so that $C^{\infty}_{0}(I_{\pm})$ may be identified with the kernel of the restriction map $C^{\infty}(S^{1})\to C_{0}^{\infty}(I_{\mp})$. Let $C_{0}^{\infty}(S^{1})$ be the space of smooth functions on $S^{1}$ that vanish to all orders at ±$\pm 1$. Let
@@
We denote by ${\mathcal{X}}$ the space of functions $F:{\mathbb{O}}\to{\mathbb{C}}$ which are smooth on ${\mathbb{O}}$, holomorphic in the interior $\mathring{{\mathbb{O}}}$, vanish at infinity, and vanish to all orders at ±$\pm 1$.
We give ${\mathcal{X}}$ the topology inherited as a subspace of $C^{\infty}(S^{1})$, that of uniform convergence of all derivatives on $S^{1}$. If $F\in{\mathcal{X}}$ we will frequently consider $F|_{I_{+}}\in C_{0}^{\infty}(I_{+})$ for the restriction to either the lower semicircle, and similarly for $I_{-}$. We will typically extend $F|_{I_{+}}$ to an element of $C^{\infty}(S^{1})$ as described above, by assigning the value 00 to the interval $I_{-}$. Thus we have
@@
By precomposing with the map $\tfrac{1}{z}$, we can consider functions holomorphic on the closed unit disk ${\mathbb{D}}=\{\left|z\right|\leq 1\}$ rather than ${\mathbb{O}}$, and without loss of generality we consider only the case of $I_{+}$. Our problem is then: fix $f\in C_{0}^{\infty}(I_{+})$, and find a sequence of functions $F_{N}$ on the closed unit disk that satisfy: (i) $F_{N}$ is smooth on ${\mathbb{D}}$ and $F_{N}(0)=0$, (ii) $F_{N}$ is holomorphic on the open unit disk $\mathring{{\mathbb{D}}}$, (iii) $F_{N}|_{I_{+}}\in C^{\infty}_{0}(I_{+})$, and (iv) $F_{N}|_{I_{+}}\to f$ in in the smooth topology. Without loss of generality we may assume that $f$ is identically zero in a neighborhood of {±$\{\pm 1\}$, as such functions are dense in $C^{\infty}_{0}(I_{+})$. We may also ignore the requirement $F_{N}(0)=0$, as if we approximate $z^{-1}f$ by $F_{N}$, then $f$ is approximated by $zF_{N}$.
@@
[…]
We have $F_{N}|_{I_{+}}\in C^{\infty}(I_{+})$ and $F_{N}|_{I_{+}}\to f$ in the smooth topology, which demonstrates points (iii) and (iv) provided we can show that $F_{N}$ extends to a (necessarily unique) function on the closed unit disk satisfying (i) and (ii). We construct the analytic extension as follows.
As before, let ${\mathbb{S}}={\mathbb{S}}_{i\pi}$ be the strip
[…]
@@
for the appropriate branch of $\log$ on the upper half-plane. This map satisfies $\alpha(\mathring{I_{+}})={\mathbb{R}}$ and $\alpha(\mathring{I_{-}})={\mathbb{R}}+i\pi$. In fact, for $z\in\mathring{I_{+}}$ we have $v_{\alpha(z)}(i)=z$, so that
[…]
Let $\tilde{f}:{\mathbb{R}}\to{\mathbb{C}}$ be the smooth function $\tilde{f}(s)=f(v_{-s}(i))$ and note that $\tilde{f}$ has compact support since $f$ was assumed to vanish in a neighborhood of ±$\pm 1$. Hence for $z\in I_{+}$ we have
[…]
$\lim_{z\to\pm 1}F_{N}(z)=0,$
and so $F_{N}$ extends to a continuous function on all of ${\mathbb{D}}$, vanishing at ±$\pm 1$.
We wish to show that the complex derivatives $F_{N}^{(k)}$, a priori defined on $\mathring{\Delta}$, also extend to continuous functions on ${\mathbb{D}}$ vanishing at ±$\pm 1$. The derivatives are given by
[…]
@@
Near $z=1$, straightforward computation yields that when $-M\leq s\leq M$ we have
$e^{-\tfrac{(N(s+\alpha(z))^{2}}{2}}\leq Ce^{-k\operatorname{Re}[\alpha(z)]^{2}}\leq Ce^{-k(\log\left|z-1\right|)^{2}}.$
As $\lim_{t\to 0^{+}}t^{-k}e^{-\log(t)^{2}}=\lim_{t\to 0^{+}}t^{-k-\log t}=0$ for any $k>0$, we see that
$\lim_{z\to 1,\left|z\right|<1}r(z)e^{-\tfrac{(N(s+\alpha(z))^{2}}{2}}=0$
and an application of the dominated convergence theorem to (3.3) yields that
$\lim_{z\to 1}F_{N}^{(k)}(z)=0.$
The same holds at $z=-1$, and we conclude that each $F_{N}^{(k)}$ extends to a continuous function on the closed disk ${\mathbb{D}}$, vanishing at ±$\pm 1$. It follows that $F_{N}$ is $C^{\infty}$ on ${\mathbb{D}}$ with the required behavior, completing the proof.
We will now study the action of analytically continued operators $\widetilde{V}_{\tau}$ on ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$, which is a dense subspace ${\mathcal{D}}$ by Lemma 3.4 and the Reeh-Schlieder property [CRTT24, App. A].
@@
The requirement that $\tau\in{\mathbb{S}}$ ensures that $v_{-\tau}(I_{+})\subset{\mathbb{O}}$, which is the domain of definition of $F$. If $f=F|_{I_{+}}\in C^{\infty}(S^{1})$ and $t\in{\mathbb{R}}$ then $\beta_{d}(v_{t})F|_{I_{+}}=\beta_{d}(v_{t})f$, which is to say that the above notation is consistent with our usual definition of $\beta_{d}(v_{t})$. Since $F$ vanishes to all orders at ±$\pm 1$ we have $\beta_{d}(v_{\tau})F|_{I_{+}}\in C_{0}^{\infty}(I_{+})$, which as previously mentioned we typically extend to a function in $C^{\infty}_{0}(S^{1})$ by extending to be the zero function on $I_{-}$. As expected we have
$\beta_{d}(v_{t})\big{[}\beta_{d}(v_{\tau})F|_{I_{+}}\big{]}=\beta_{d}(v_{\tau+t})F|_{I_{+}}.$ (3.5)
@@
$\widetilde{V}_{i\pi}\Phi=(-1)^{d+k}\varphi_{1}(F_{1}\circ z^{-1}|_{I_{+}})\cdots\varphi_{k}(F_{k}\circ z^{-1}|_{I_{+}})\Omega.$ (3.8)
@@
Indeed $F\circ z^{-1}$ is holomorphic on the disk $\mathring{{\mathbb{D}}}$ and vanishes at the origin, and thus $\varphi(F\circ z^{-1}|_{S^{1}})$ strictly lowers conformal dimension. In particular,

### v2

@@ 2.2 Wightman CFTs and vertex algebras
Our main object of study is potentially non-unitary Möbius-covariant Wightman conformal field theories on the unit circle $S^{1}\subset{\mathbb{C}}$. In this context, the Möbius group $\operatorname{M\ddot{o}b}$ refers to $\operatorname{PSU}(1,1)$, the group of holomorphic automorphisms of the unit disk. Such theories were systematically studied in [CRTT25] (see also [RTT22]). In contrast to Section 2.1, we do not limit ourselves to unitary theories (which possess a compatible inner product). We will summarize here the basic notions, and the reader may consult the reference for further details.
@@
We use the Fréchet topology induced on $C^{\infty}(S^{1})$ by the $C^{N}$ norms [^4] In [RTT22, CRTT25] we used the Sobolev norms $\left\|f\right\|_{H^{N}}$, but the resulting topologies are equivalent due to embeddings $H^{N}(S^{1})\subset C^{N^{\prime}}(S^{1})$ and $C^{M}(S^{1})\subset H^{M^{\prime}}(S^{1})$.
$\left\|f\right\|_{C^{N}}=\sum_{j=0}^{N}\big\|f^{(j)}\big\|_{\infty}=\sum_{j=0}^{N}\big\|\tfrac{d^{j}}{d\theta^{j}}f\big\|_{\infty}.$
@@
The Wightman axioms have not been extensively studied in their own right within the context of axiomatic 2d chiral conformal field theory, but it turns out that they are exactly equivalent to Möbius vertex algebras, along with a chosen set of quasiprimary vectors which generate the vertex algebra. It has been long-known that this correspondence is “morally valid” (see e.g. [Kac98, §1.2]), certainly in the case of unitary theories, but it was made precise in [CRTT25]. In the case of vertex algebras, the Möbius covariance condition is with respect to the complexified Lie algebra $\operatorname{Lie}(\operatorname{M\ddot{o}b})_{\mathbb{C}}$, which is spanned by elements $\{L_{-1},L_{0},L_{1}\}$ satisfying the familiar relation $L_{n},L_{m}]=(n-m)L_{n+m}$.
@@
    $\Omega$ is $\operatorname{Lie}(\operatorname{M\ddot{o}b})$-invariant, i.e. $L_{m}\Omega=0$ for m=1,0,1m=-1,0,1.
@@
    $L_{m},Y(v,z)]=\sum_{j=0}^{m+1}\binom{m+1}{j}z^{m+1-j}Y(L_{j-1}v,z)$ and $Y(L_{-1}v,z)=\frac{d}{dz}Y(v,z)$ for all $v\in{\mathcal{V}}$ and m=1,0,1m=-1,0,1.
@@
In the other direction, given a Möbius vertex algebra $\mathcal{V}$} with a generating set of quasiprimary vectors, there is a canonical extension ${\mathcal{D}}\supset{\mathcal{V}}$ (which can be realized as a subspace of the algebraic completion of $\mathcal{V}$}) and Möbius-covariant Wightman CFT on $\mathcal{D}$} corresponding to $\mathcal{V}$} in the sense described above. This correspondence can be made into an equivalence of categories [CRTT25, §3.3].
@@
In both settings, we have a notion of an invariant Hermitian form, which we call an involutive theory in general, and a unitary theory if the Hermitian form is an inner product. Some key examples of involutive but non-unitary theories include the non-unitary minimal models, such as the Yang-Lee model. These notions will be outlined in detail in Sections 4 and 6, respectively. Invariant forms for a Wightman theory exactly correspond to invariant forms for the corresponding vertex algebra [CRTT25, §4].
We close with a brief preview of the themes of the article. Following [SW64], given a Wightman CFT $\mathcal{F}$} acting on its domain $\mathcal{D}$}, for every interval [^5] The term ‘interval’ in $S^{1}$ will be used exclusively for open, nonempty, proper, non-dense connected subsets of $S^{1}$. $I\subset S^{1}$ we form the polynomial algebra ${\mathcal{P}}(I)\subset\operatorname{End}({\mathcal{D}})$, which is to say the unital algebra generated by operators $\varphi(f)$ with $\operatorname{supp}(f)\subset I$. We obtain a net $\mathcal{P}$} of algebras which formally satisfies many of the properties of a Haag-Kastler net, but is in fact encoding the data of a Möbius vertex algebra. From [CRTT25, App. A], this net has the Reeh-Schlieder property, which is to say that ${\mathcal{P}}(I)\Omega$ is ($\mathcal{F}$}-strong) dense in $\mathcal{D}$} for every interval $I$, and if $x\in{\mathcal{P}}(I)$ and $x\Omega=0$, then $x=0$. Continuing along the same line, we have the following motivating questions:
@@ 3 Analytic continuation for non-unitary Wightman CFTs
For an interval $I\subset S^{1}$ we let ${\mathcal{P}}(I)\subset\operatorname{End}({\mathcal{D}})$ be the unital algebra generated by operators $\varphi(f)$ with $\varphi\in{\mathcal{F}}$ and $\operatorname{supp}f\subset I$. For $\tau\in{\mathbb{C}}$, let ${\mathbb{S}}_{\tau}$ be the closed strip bounded by the lines $\mathbb{R}$} and ${\mathbb{R}}+\tau$. Our first goal is to investigate a certain weak form of analytic continuation of $V_{t}$ to the closed strip
${\mathbb{S}}:={\mathbb{S}}_{i\pi}=\{\tau\in{\mathbb{C}}\,:\,0\leq\operatorname{Im}(\tau)\leq\pi\}.$
@@
are (jointly) continuous in the $f_{j}\in C^{\infty}(S^{1})$, for every $\Phi\in{\mathcal{D}}$.
@@
Note that such a vector $\Psi$ is uniquely determined by $\Phi$ (i.e. $\widetilde{V}_{\tau}$ is well-defined) by the regularity property of the Wightman CFT, which asserts that $\mathcal{D}$$_{\mathcal{F}}^{*}$ separates points.
If $\tau=t\in{\mathbb{R}}$ then $D(\widetilde{V}_{t})={\mathcal{D}}$ and $\widetilde{V}_{t}=V_{t}$. The tilde in the notation $\widetilde{V}_{\tau}$ and $D(\widetilde{V}_{\tau})$ is present to distinguish from the domain $D(V_{\tau})$ considered in Section 5 in the context of CFTs with an invariant Hermitian form. If $\tau\in{\mathbb{C}}$ and $t\in{\mathbb{R}}$ we have $\widetilde{V}_{\tau+t}=\widetilde{V}_{\tau}V_{t}=V_{t}\widetilde{V}_{\tau}$ as partially defined operators [^7] This means that the operators in question have the same domain, where the domain of the composite of partially defined operators $X_{1}X_{2}$ consists of all vectors $\Phi$ such that $\Phi\in D(X_{2})$ and $X_{2}\Phi\in D(X_{1})$.. More generally for $\tau_{1},\tau_{2}\in{\mathbb{C}}$ the operator $\widetilde{V}_{\tau_{1}+\tau_{2}}$ is an extension of $\widetilde{V}_{\tau_{1}}\widetilde{V}_{\tau_{2}}$. It is clear that $\Omega\in D(\widetilde{V}_{\tau})$, but it is not immediately obvious that $D(\widetilde{V}_{\tau})$ is any larger. Thus our first step is to identify a ‘large’ subspace of $D(\widetilde{V}_{\tau})$.
Let $C_{0}^{\infty}(I_{+})$ and $C_{0}^{\infty}(I_{-})$ be smooth functions on $I_{+}$ (resp. $I_{-}$) that vanish to all orders at the endpoints. We will sometimes embed these spaces into $C^{\infty}(S^{1})$ by extending with the value 0, so that $C^{\infty}_{0}(I_{\pm})$ may be identified with the kernel of the restriction map $C_{0}^{\infty}(S^{1})\to C_{0}^{\infty}(I_{\mp})$, where $C_{0}^{\infty}(S^{1})$ is the space of smooth functions on $S^{1}$ that vanish to all orders at $\pm 1$. Let
@@
We denote by $\mathcal{X}$} the space of functions $F:{\mathbb{O}}\to{\mathbb{C}}$ which are smooth on $\mathbb{O}$}, holomorphic in the interior $\mathring{{\mathbb{O}}}$, vanish at infinity, and vanish to all orders at $\pm 1$.
We give $\mathcal{X}$} the topology inherited from its embedding into $C^{\infty}(S^{1})$ via restriction, namely the topology of uniform convergence of all derivatives on $S^{1}$, which we call the $C^{\infty}$ topology. If $F\in{\mathcal{X}}$ we will frequently consider $F|_{I_{+}}\in C_{0}^{\infty}(I_{+})$ for the restriction to the upper semicircle, and similarly for $I_{-}$. We will typically extend $F|_{I_{+}}$ to an element of $C^{\infty}(S^{1})$ as described above, by assigning the value 0 to the interval $I_{-}$. Thus we have
@@
By precomposing with the map $\tfrac{1}{z}$, we can consider functions holomorphic on the closed unit disk ${\mathbb{D}}=\{\left|z\right|\leq 1\}$ rather than $\mathbb{O}$}, and without loss of generality we consider only the case of $I_{+}$. Our problem is then: fix $f\in C_{0}^{\infty}(I_{+})$, and find a sequence of functions $F_{N}$ on the closed unit disk that satisfy: (i) $F_{N}$ is smooth on $\mathbb{D}$} and $F_{N}(0)=0$, (ii) $F_{N}$ is holomorphic on the open unit disk $\mathring{{\mathbb{D}}}$, (iii) $F_{N}|_{I_{+}}\in C^{\infty}_{0}(I_{+})$, and (iv) $F_{N}|_{I_{+}}\to f$ in the $C^{\infty}$ topology. Without loss of generality we may assume that $f$ is identically zero in a neighborhood of $\{\pm 1\}$, as such functions are dense in $C^{\infty}_{0}(I_{+})$. We may also ignore the requirement $F_{N}(0)=0$, as if we approximate $z^{-1}f$ by $F_{N}$, then $f$ is approximated by $zF_{N}$.
@@
$F_{N}|_{I_{+}}:=\frac{N}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{-\tfrac{(Ns)^{2}}{2}}\cdot f\circ v_{-s}\,ds.$ (3.1)
We have $F_{N}|_{I_{+}}\in C^{\infty}(I_{+})$ and $F_{N}|_{I_{+}}\to f$ in the $C^{\infty}$ topology, which demonstrates points (iii) and (iv) provided we can show that $F_{N}$ extends to a (necessarily unique) function on the closed unit disk satisfying (i) and (ii). We will do this in two steps. In Step 1, we will show that $F_{N}$ extends from its original domain $I_{+}$ to a function holomorphic in a neighborhood of ${\mathbb{D}}\setminus\{\pm 1\}$. In Step 2, we will show that $F_{N}$ extends further to take the value zero at $z=\pm 1$, and that the resulting function is smooth on the closed disk $\mathbb{D}$}. These two steps demonstrate the existence of the required sequence $F_{N}$, at which point the proof will be complete.
We begin with Step 1, and construct an analytic extension of $F_{N}$ to a neighborhood of ${\mathbb{D}}\setminus\{\pm 1\}$ as follows. As before, let ${\mathbb{S}}={\mathbb{S}}_{i\pi}$ be the strip
${\mathbb{S}}=\{z\in{\mathbb{C}}\,:\,0\leq\operatorname{Im}(z)\leq\pi\}$
@@
with the branch of $\log$ on the upper half-plane chosen so that the image of $\alpha$ is $\mathbb{S}$}. This map satisfies $\alpha(\mathring{I_{+}})={\mathbb{R}}$ and $\alpha(\mathring{I_{-}})={\mathbb{R}}+i\pi$. In fact, for $z\in\mathring{I_{+}}$ we have $v_{\alpha(z)}(i)=z$, so that
$F_{N}(z)$ $=\frac{N}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{-\tfrac{(Ns)^{2}}{2}}\,f(v_{-s}(z))\,ds$
$=\frac{N}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{-\tfrac{(Ns)^{2}}{2}}\,f(v_{-s+\alpha(z)}(i))\,ds$
$=\frac{N}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{-\tfrac{(N(s+\alpha(z)))^{2}}{2}}\,f(v_{-s}(i))\,ds.$
Let $\tilde{f}$$\mathbb{R}$$\to{\mathbb{C}}$ be the smooth function $\tilde{f}(s)=f(v_{-s}(i))$ and note that $\tilde{f}$ has compact support since $f$ was assumed to vanish in a neighborhood of $\pm 1$. Hence for $z\in\mathring{I}_{+}$ we have
$F_{N}(z)=\frac{N}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{-\tfrac{(N(s+\alpha(z)))^{2}}{2}}\,\tilde{f}(s)\,ds=\frac{N}{\sqrt{2\pi}}\int_{-M}^{M}e^{-\tfrac{(N(s+\alpha(z)))^{2}}{2}}\,\tilde{f}(s)\,ds,$ (3.2)
for some $M$ sufficiently large that the interval [M,M][-M,M] contains the support of $\tilde{f}$. The formula (3.2) now visibly defines a holomorphic function of $z$ in a neighborhood of ${\mathbb{D}}\setminus\{\pm 1\}$ which extends the original definition of $F_{N}$ on $\mathring{I}_{+}$, and we denote this extension again by $F_{N}$. This completes Step 1.
In Step 2, we will show that $F_{N}$ extends further to a smooth function on the closed disk $\mathbb{D}$}. We do this by showing that the holomorphic function $F_{N}|_{\mathring{{\mathbb{D}}}}$ and all of its derivatives extend continuously to $\partial{\mathbb{D}}=S^{1}$. From Step 1 we know that there exists such an extension to a neighborhood of ${\mathbb{D}}\setminus\{\pm 1\}$, so it remains to consider the behavior near $z=\pm 1$. We first argue that the integrand of (3.2) is bounded as a function of $(s,z)\in[-M,M]\times({\mathbb{D}}\setminus\{\pm 1\})$. By inspection, it suffices to show that the function $\operatorname{Re}\big[(s+\alpha(z))^{2}\big]$ is bounded below on the same domain. And indeed,
$\operatorname{Re}\big[(s+\alpha(z))^{2}\big]=-\operatorname{Im}[\alpha(z)]^{2}+(s+\operatorname{Re}[\alpha(z)])^{2}$
is visibly bounded below, since $\alpha$ takes values in the strip $\mathbb{S}$}.
Since the integrand of (3.2) is bounded, we may apply the dominated convergence theorem to obtain
$\lim_{z\to\pm 1}F_{N}(z)=\frac{N}{\sqrt{2\pi}}\int_{-M}^{M}\lim_{z\to\pm 1}e^{-\tfrac{(N(s+\alpha(z)))^{2}}{2}}\,\tilde{f}(s)\,ds=0,$
with the limit taken over $z\in{\mathbb{D}}\setminus\{\pm 1\}$. It follows that $F_{N}$ extends to a continuous function on all of $\mathbb{D}$}, vanishing at $\pm 1$.
We next wish to show that the complex derivatives $F_{N}^{(k)}$, a priori defined on $\mathring{{\mathbb{D}}}$, also extend to continuous functions on $\mathbb{D}$} vanishing at $\pm 1$. We use a similar argument. Since the integrand of (3.2) is smooth on $M,M]\times({\mathbb{D}}\setminus\{\pm 1\})$, we may differentiate under the integral and see that the iterated $z$-derivatives of $F_{N}$ are of the form
$F^{(k)}_{N}(z)=\frac{N}{\sqrt{2\pi}}\int_{-M}^{M}r(z)e^{-\tfrac{(N(s+\alpha(z)))^{2}}{2}}\,\tilde{f}(s)\,ds,$ (3.3)
@@
Near $z=1$, straightforward computation yields that when $M\leq s\leq M$ we have
$e^{-\tfrac{(N(s+\alpha(z)))^{2}}{2}}\leq Ce^{-k\operatorname{Re}[\alpha(z)]^{2}}\leq Ce^{-k(\log\left|z-1\right|)^{2}}.$
As $\lim_{t\to 0^{+}}t^{-k}e^{-(\log t)^{2}}=\lim_{t\to 0^{+}}t^{-k-\log t}=0$ for any $k>0$, we see that
$\lim_{z\to 1}r(z)e^{-\tfrac{(N(s+\alpha(z)))^{2}}{2}}=0,$
with the limit taken over $z\in{\mathbb{D}}\setminus\{\pm 1\}$. An application of the dominated convergence theorem to (3.3) now yields that
$\lim_{z\to 1}F_{N}^{(k)}(z)=0,$
with the limit again taken over $z\in{\mathbb{D}}\setminus\{\pm 1\}$. The same holds at z=1z=-1, and we conclude that each $F_{N}^{(k)}$ extends to a continuous function on the closed disk $\mathbb{D}$}, vanishing at $\pm 1$. It follows that $F_{N}$ extends to a smooth function on $\mathbb{D}$} with the required behavior, completing the proof.
We will now study the action of analytically continued operators $\widetilde{V}_{\tau}$ on ${\mathcal{P}}_{\mathcal{X}}(I_{+})\Omega$, which is a dense subspace $\mathcal{D}$} by Lemma 3.4 and the Reeh-Schlieder property [CRTT25, App. A].
@@
The requirement that $\tau\in{\mathbb{S}}$ ensures that $v_{-\tau}(I_{+})\subset{\mathbb{O}}$, which is the domain of definition of $F$. If $f=F|_{I_{+}}\in C^{\infty}(S^{1})$ and $t\in{\mathbb{R}}$ then $\beta_{d}(v_{t})F|_{I_{+}}=\beta_{d}(v_{t})f$, which is to say that the above notation is consistent with our usual definition of $\beta_{d}(v_{t})$. Since $F$ vanishes to all orders at $\pm 1$ we have $\beta_{d}(v_{\tau})F|_{I_{+}}\in C_{0}^{\infty}(I_{+})$, which as previously mentioned we typically extend to a function in $C^{\infty}_{0}(S^{1})$ by extending to be the zero function on $I_{-}$. It follows from the definition that
$\beta_{d}(v_{t})\big[\beta_{d}(v_{\tau})F|_{I_{+}}\big]=\beta_{d}(v_{\tau+t})F|_{I_{+}},$ (3.5)
giving an analytic continuation of the more standard identity that holds when $\tau\in{\mathbb{R}}$.
@@
$\widetilde{V}_{i\pi}\Phi=(-1)^{d-k}\varphi_{1}(F_{1}\circ z^{-1}|_{I_{+}})\cdots\varphi_{k}(F_{k}\circ z^{-1}|_{I_{+}})\Omega.$ (3.8)
@@
Indeed $F\circ z^{-1}$ is holomorphic on the disk $\mathring{{\mathbb{D}}}$ and vanishes at the origin, and hence its power series expansion at the origin contains only (strictly) positive powers of $z$. It follows that $\varphi(F\circ z^{-1}|_{S^{1}})$ strictly lowers conformal dimension. In particular,